# Non-Uniqueness of Metasurfaces for Wave Transformations

### v1

@@ Abstract
We show that a large family of tensorial metasurfaces can be found that perform an identical wave transformation, showing that even when the conditions of reciprocity and passivity are imposed, there still remain many solutions to the design problem. As an example, we explore the case of a metasurface that rotates a single input polarization, showing we can parameterize the set of equivalent reciprocal metasurfaces in terms of a single complex parameter. Through allowing dissipation and gain within the response, the surface can have many different functionalities in the orthogonal polarization, opening up a new route for the design of multiplexed metasurfaces.
@@ I Introduction
Metasurfaces are planar media containing an array of polarizable elements Chen et al. (2016), carefully chosen to perform a desired transformation of an incoming wave. Although this concept covers many areas of wave physics Assouar et al. (2018), here the focus is on metasurfaces for electromagnetic waves. Within such surfaces, the elements are typically spaced closer than the wavelength, allowing the characterization of the array in terms of an effective electromagnetic boundary condition Achouri and Caloz (2021). Whereas partially or totally transmitting metasurfaces are typically characterized in terms of sheet transition conditions Tretyakov (2003); Wu et al. (2018); Lebbe et al. (2023), or a polarizability Holloway et al. (2009); Zalus̆ki et al. (2016), impenetrable surfaces are usually characterized in terms of a surface impedance Senior and Volakis (1995); Tretyakov (2003); Maci et al. (2011). Illustrated in Fig. 1. This amounts to specifying the ratio of the electric and magnetic fields at all points on the surface. There are a variety of approaches to the design of such metasurfaces, controlling e.g. the local reflection phase over the surface Yu and Capasso (2014); Yu et al. (2011a); Chen et al. (2014a); Wang et al. (2019), or the coupling of surface waves to free–space radiation Fong et al. (2010). In this work, we consider impenetrable metasurfaces described in terms of a tensorial surface impedance and investigate the uniqueness of the surface parameters with respect to a desired transformation of the field.
@@
Figure 1: Field transformations with tensorial metasurfaces: A metasurface characterized in terms of a tensorial surface impedance, $\mathrm{Z}_{s}$ can perform an arbitrary transformation of an incident electromagnetic wave. Here a normally incident wave with electric field aligned along the $x$–axis is transformed into a normally reflected wave with electric field at an angle $\theta$ to the $x$–axis, an example treated in Sec. IV. The right hand panels (a,b) illustrate the behaviour of a tensorial metasurface in terms of the induced current $\mathbf{J}_{\parallel}$ and the electric field eigenvectors, $\mathbf{E}_{\parallel,i}$, in terms of which (ci,cii) any incident field can be decomposed
@@
Although the terminology of a ‘metasurface’ is little more than a decade old, the idea is a continuation of a large body of earlier work on frequency selective surfaces Munk (2005), which was revived when similar ideas began to be implemented in optics Yu et al. (2011b). Recent work has taken the metasurface idea significantly further forward. For example, through matching the electric and magnetic polarizability of the surface elements in a transmitting sheet, all reflection can be eliminated, resulting in so–called ‘Huygens’ metasurfaces Pfeiffer and Grbic (2013); Epstein and Eleftheriades (2016); Chen et al. (2018), which sculpt the field while maintaining its direction of propagation. More recently, more exotic metasurfaces have been considered, including non–Hermitian and PT–symmetric elements Fleury et al. (2014); Tapar and Kishen (2021); Y. et al. (2022), which contain carefully designed distributions of loss and/or gain, allowing for an unusual control over e.g. the polarization of the output field. Novel forms of surface excitations have also been discovered at junctions between metasurfaces Horsley and Hooper (2014); Bisharat and Sievenpiper (2017), and most recently non–reciprocal Lawrence et al. (2018) and space–time varying surface parameters have been considered Li et al. (2020); Wang et al. (2020), enabling both multiplexed functionality and spectral control.
@@
As mentioned above, here we investigate the uniqueness of the tensorial surface impedance, $\mathrm{Z}_{s}$, when enforcing a single given transformation of the electromagnetic field. This is relevant to the problem of designing multiplexed—more specifically polarization multiplexed—metasurfaces, where two or more functionalities are imposed in the same design. Unless the surface properties are e.g. re–configurable Shirmanesh et al. (2020), it is well known that this is a difficult design problem, and there have been numerous previous approaches Capers et al. (2022); Dai et al. (2023); Murtaza Rind et al. (2024), some of which are summarized in Ref. Xu et al. (2021).
Here we show that, once a single field transformation has been specified at every point on the surface then there remains a great deal of freedom in the parameters determining the surface impedance. The greatest freedom occurs for non–reciprocal surfaces, but even reciprocal ones always contain a single free complex parameter. When we constrain the surface to be both passive and reciprocal, we find that through increasing dissipation within the surface, and thereby decreasing the amplitude of the transformed field, we, in a sense, increase the non–uniqueness of the metasurface. This shows that, through carefully increasing dissipation within the structure, we gain more freedom to find a metasurface that has a specific polarization multiplexed function.
@@ II The Tensorial Impedance Boundary Condition
We can characterize the reflection of electromagnetic waves from any thin impenetrable metasurface in terms of an impedance boundary condition written in terms of a $2\times 2$ tensor valued surface impedance Senior and Volakis (1995); Tretyakov (2003), $\mathrm{Z}_{s}$. To understand the origin of this boundary condition, consider Ohm’s law within the metasurface relating the electric field parallel to a boundary, $\mathbf{E}_{\parallel}$, to the surface currents, $\mathbf{j}_{s}$, via the surface conductivity, $\sigma_{s}$. By relating the surface currents to the magnetic field via $\mathbf{j}_{s}=\hat{\mathbf{n}}\times\mathbf{H}_{\parallel}$ where $\hat{\mathbf{n}}$ is the surface normal, we have an equation relating the surface electric and magnetic fields,
@@
$\mathbf{E}_{\parallel}=\mathrm{Z}_{s}\cdot\left(\hat{\mathbf{n}}\times\mathbf{H}_{\parallel}\right).$ (2)
@@
Given our assumption that the metasurface is impenetrable, a positive value for $\langle P_{n}\rangle$ indicates a dissipative response, whereas a negative value indicates amplification, or gain. Given that a lossless metasurface will have $\langle P_{n}\rangle=0$ for all possible incident fields, lossless surfaces have an impedance tensor that is ‘${\rm i}$’ times a Hermitian matrix (a so–called anti–Hermitian operator). In the general case, we can decompose the impedance into Hermitian and anti–Hermitian parts, $Z_{s}=H_{1}+{\rm i}\,H_{2}$, where $H_{1,2}$ are Hermitian matrices. The dissipated power is thus proportional to the Rayleigh quotient of $H_{1}$, implying that passive surfaces (i.e. those without gain) have a positive semi–definite form of $H_{1}$.
@@
III Field transformations with tensorial metasurfaces
An important application of these metasurfaces is in wave transformations, where a given incident wave is reflected into a specified radiation pattern, which in general can include both near and far–fields. Such transformations include the redirection, focusing, and absorption of an incident wave. One simple method for finding the impedance distribution that enacts a given transformation is to directly invert Eq. (2). To do this, we write our complex valued surface fields in terms of their magnitude, e.g. $|E|=(\mathbf{E}_{\parallel}\cdot\mathbf{E}_{\parallel}^{\star})^{1/2}$ and “direction” $\hat{\mathbf{e}}=\mathbf{E}_{\parallel}/|E|$,
$\mathbf{E}_{\parallel}$ $=|E|\hat{\mathbf{e}}$
$\mathbf{H}_{\parallel}$ $=|H|\hat{\mathbf{h}}$ (4)
Substituting this decomposition into the earlier expression for the impedance boundary condition (2) we obtain the requirement
@@ III Field transformations with tensorial metasurfaces
which tells us that at each point on the surface, the impedance tensor must rotate and rescale the complex unit vector $\hat{\mathbf{n}}\times\hat{\mathbf{h}}$ so that it is parallel to the electric field unit vector, $\hat{\mathbf{e}}$. When our desired field transformation has $\hat{\mathbf{e}}$ everywhere parallel to $\hat{\mathbf{n}}\times\hat{\mathbf{h}}_{\parallel}$, this can be achieved with an isotropic metasurface, characterized in terms of a scalar impedance, $\mathrm{Z}_{s}=|E|/|H|$. This is possible in e.g. a $2D$ geometry, where one field component always lies out of the plane and the other always in the plane.
Nevertheless, in the general case a tensor valued impedance is required to satisfy (5). A simple solution is $\mathrm{Z}_{s}=(|E|/|H|)\,\hat{\mathbf{e}}\otimes(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}$, where $\otimes$’ is the tensor product defined as e.g. $(\mathbf{a}\otimes\mathbf{b})_{ij}=\left(\mathbf{a}\mathbf{b}^{T}\right)_{ij}=a_{i}b_{j}$. However, as also explained in Ref. Kwon (2020), this form of surface impedance is most often neither passive ($\mathrm{Z}_{s}+\mathrm{Z}_{s}^{\dagger}$ not positive semi–definite), nor reciprocal ($\mathrm{Z}_{s}\neq\mathrm{Z}_{s}^{\rm T}$). This certainly represents a challenge for any practical realizations of these metasurface parameters, as both a power source and e.g. magnetic elements would have to be embedded within the surface! Yet there is an additional freedom that we can exploit to bring the design closer to reality. We can always perform a transformation $\mathrm{Z}_{s}\to\mathrm{Z}_{s}+\mathbf{v}\otimes\hat{\mathbf{h}}$, where $\mathbf{v}$ is an arbitrary vector, without affecting condition (5). This leaves us with an impedance tensor of the general form,
$\mathrm{Z}_{s}=\frac{|E|}{|H|}\left[\hat{\mathbf{e}}\otimes\left(\hat{\mathbf{n}}\times\hat{\mathbf{h}}\right)^{\star}+\mathbf{v}\otimes\hat{\mathbf{h}}\right].$ (6)
This surface impedance performs the desired transformation (5). Generally, it is non–reciprocal and non–passive, with the freedom to choose both complex components of $\mathbf{v}$, arbitrarily. The remainder of this paper explores this freedom in the design of metasurfaces.
@@ III.1 Reciprocal metasurfaces:
To illustrate special cases of Eq. (6), we adopt a complex basis to express the impedance tensor. Here, we use the complex basis vectors $\hat{\mathbf{h}}$ and $\hat{\mathbf{n}}\times\hat{\mathbf{h}}^{\star}$ and the dual basis, $\hat{\mathbf{h}}^{\star}$ and $\hat{\mathbf{n}}\times\hat{\mathbf{h}}$, with the electric field direction being given by $\hat{\mathbf{e}}=e_{1}\,\hat{\mathbf{h}}+e_{2}\,(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}=\bar{e}_{1}\,\hat{\mathbf{h}}^{\star}+\bar{e}_{2}\,(\hat{\mathbf{n}}\times\hat{\mathbf{h}})$ and the arbitrary vector is written as $\mathbf{v}=\alpha\,\hat{\mathbf{h}}+\beta\,(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}=\bar{\alpha}\hat{\mathbf{h}}^{\star}+\bar{\beta}(\mathbf{n}\times\hat{\mathbf{h}})$. In terms of these basis vectors the impedance tensor (6) becomes
$\mathrm{Z}_{s}=\frac{|E|}{|H|}\bigg{\{}e_{1}\left[\,\hat{\mathbf{h}}\otimes(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}+\,(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}\otimes\hat{\mathbf{h}}\right]+e_{2}\,(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}\otimes(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}+\alpha\,\hat{\mathbf{h}}\otimes\hat{\mathbf{h}}\bigg{\}}$ (7)
@@
From Eq. (7), we can see that, irrespective of the field transformation we seek to enact, it is always possible to find a reciprocal metasurface that will perform this transformation. Indeed, we have an infinite family of such surfaces as we are still free to choose the complex valued parameter $\alpha$. Provided we can implement amplification across the metasurface, this shows that there is a non-uniqueness in the surface parameters. The physical reason for this design freedom is that at every point we are imposing the response to a given surface field. Of course this leaves the response to the orthogonal surface field unspecified. The somewhat surprising fact is that while the requirement of reciprocity removes some of this redundancy, it is not enough to completely determine the form of the surface impedance.
@@ III.2 Passive metasurfaces:
This is a requirement on the field transformation that we must implement at every point on the surface at the design stage: we cannot hope for a passive surface impedance if our design fields require the metasurface to generate energy! However, even assuming Eq. (9) holds, it may still be that the impedance (7) exhibits gain when probed with other incident fields (III).
@@
Fig. 2 uses a randomly generated set of surface electric and magnetic fields to show that imposing this constraint indeed results in a passive metasurface. Note that if we impose zero dissipation in the response to the design field, ${\rm Re}[\bar{e}_{2}]=0$, the inequality (12) is impossible to fulfill. This shows that reducing the efficiency of the design (i.e. decreasing the reflected power) can help to find a set of metasurface parameters that are, overall passive. We shall return to this point again in the next section.
@@
Figure 2: Randomly generated wave transformations: Using a set of $500$ randomly generated complex unit vectors, $\hat{\mathbf{e}}$ and $\hat{\mathbf{h}}$, representing different choices of wave transformation, we computed the reciprocal surface impedance given by Eq. (7) (with $|E|/|H|=1$, plus the constraint (9)), and the eigenvalues of $(\mathrm{Z}_{s}+\mathrm{Z}_{s}^{\dagger})/2$, which determine the dissipation/gain in the metasurface. (a) $\alpha=0$ for all designs, giving rise to eigenvalues of both signs, generally requiring a non–passive metasurface. (b) $\alpha$ constrained by Eq. (12) (here taking the equality), giving rise to eigenvalues of a positive sign only, showing that all surfaces can be made passive, as desired. (c) Corresponding values of $\alpha$ for the metasurfaces considered in panel (b), showing that in general this parameter must be complex valued.

### v3

# Finding passive, reciprocal metasurfaces for arbitrary wave transformations

@@ Abstract
We give a general design method for finding the passive, reciprocal surface impedance tensor required to enact any wave transformation. We do this through characterising the surface in terms of a tensorial surface impedance, showing that a large family of impedance distributions can be found that perform an identical wave transformation. Even when the conditions of reciprocity and passivity are imposed, there still remain many solutions to the design problem. We exploit this as a design method for metasurfaces, giving two examples where the metasurface rotates the input polarization and reshapes the output field, showing we can parameterize the set of equivalent reciprocal metasurfaces in terms of a single complex parameter. In addition, through allowing dissipation and gain within the response, the surface can have many different functionalities in the orthogonal polarization, opening up a new route for the design of multiplexed metasurfaces.
@@ I Introduction
Metasurfaces are planar media containing an array of polarizable elements Chen et al. (2016),carefully chosen to perform a desired transformation of an incoming wave. Although this concept covers many areas of wave physics Assouar et al. (2018),here we focus on metasurfaces for electromagnetic waves. Within such surfaces, the elements are typically spaced closer than the wavelength, allowing the characterization of the array in terms of an effective electromagnetic boundary condition Achouri and Caloz (2021). Whereas partially or totally transmitting metasurfaces are typically characterized in terms of sheet transition conditions Tretyakov (2003); Wu et al. (2018); Lebbe et al. (2023), or a polarizability Holloway et al. (2009); Zalus̆ki et al. (2016), impenetrable surfaces are usually characterized in terms of a surface impedance Senior and Volakis (1995); Tretyakov (2003); Maci et al. (2011), as illustrated in Fig. 1. This amounts to specifying the ratio of the electric and magnetic fields at all points on the surface. There are a variety of approaches to the design of such metasurfaces, controlling e.g. the local reflection phase over the surface Yu and Capasso (2014); Yu et al. (2011); Chen et al. (2014a); Wang et al. (2019), or the coupling of surface waves to free–space radiation Fong et al. (2010). In this work, we consider impenetrable metasurfaces described in terms of a tensorial surface impedance. As we shall show, there is a non–uniqueness of the surface parameters with respect to a desired transformation of the field. Besides being generally interesting, this can be exploited to impose the practical conditions of passivity and reciprocity on any metasurface design. The greatest freedom occurs for non–reciprocal surfaces, but even reciprocal ones always contain a single free complex parameter.
@@
Figure 1: Field transformations with tensorial metasurfaces: A metasurface characterized in terms of a tensorial surface impedance, $\mathrm{Z}_{s}$ can perform an arbitrary transformation of an incident electromagnetic wave. Here a normally incident wave with electric field aligned along the xx–axis is transformed into a normally reflected wave with electric field at an angle $\theta$ to the xx–axis, an example treated in Sec. IV. The right hand panels (a,b) illustrate the behaviour of a tensorial metasurface in terms of the induced current $\mathbf{J}_{\parallel}$ and the electric field eigenvectors, $\mathbf{E}_{\parallel,i}$, in terms of which (ci,cii) any incident field can be decomposed
@@
Although the terminology of a ‘metasurface’ is little more than a decade old, the idea is a continuation of a large body of earlier work on frequency selective surfaces Munk (2005), which was revived when similar ideas began to be implemented in optics Yu et al. (2011). Recent work has taken the metasurface idea significantly further forward. For example, through matching the electric and magnetic polarizability of the surface elements in a transmitting sheet, all reflection can be eliminated, resulting in so–called ‘Huygens’ metasurfaces Pfeiffer and Grbic (2013); Epstein and Eleftheriades (2016); Chen et al. (2018), which sculpt the field while maintaining its direction of propagation. More recently, more exotic metasurfaces have been considered, including non–Hermitian and PT–symmetric elements Fleury et al. (2014); Tapar and Kishen (2021); Y. et al. (2022), which contain carefully designed distributions of loss and/or gain, allowing for an unusual control over e.g. the polarization of the output field. Novel forms of surface excitations have also been discovered at junctions between metasurfaces Horsley and Hooper (2014); Bisharat and Sievenpiper (2017), and most recently non–reciprocal Lawrence et al. (2018) and space–time varying surface parameters have been considered Li et al. (2020); Wang et al. (2020), enabling both multiplexed functionality and spectral control.
@@
As mentioned above, here we investigate the uniqueness of the tensorial surface impedance, $\mathrm{Z}_{s}$, when enforcing a single given transformation of the electromagnetic field. As well as exploiting this to find practical (reciprocal and passive) sets of parameters, this finding is also relevant to the problem of designing multiplexed—more specifically polarization multiplexed—metasurfaces, where two or more functionalities are imposed in the same design. Unless the surface properties are e.g. re–configurable Shirmanesh et al. (2020), it is well known that this is a difficult design problem, and there have been numerous previous approaches Capers et al. (2022); Dai et al. (2023); Murtaza Rind et al. (2024), some of which are summarized in Ref. Xu et al. (2021).
@@ II The Tensorial Impedance Boundary Condition
In this section we review the mathematical description of impenetrable metasurfaces in terms of the impedance boundary condition. We can characterize the reflection of electromagnetic waves from any thin impenetrable metasurface in terms of an impedance boundary condition written in terms of a $2\times 2$ tensor valued surface impedance Senior and Volakis (1995); Tretyakov (2003), $\mathrm{Z}_{s}$. To understand the origin of this boundary condition, consider Ohm’s law within the metasurface relating the electric field parallel to a boundary, $\mathbf{E}_{\parallel}$, to the surface currents, $\mathbf{j}_{s}$, via the surface conductivity, $\sigma_{s}$. By relating the surface currents to the magnetic field via $\mathbf{j}_{s}=\hat{\mathbf{n}}\times\mathbf{H}_{\parallel}$ where $\hat{\mathbf{n}}$ is the surface normal, we have an equation relating the surface electric and magnetic fields,
@@
$\mathbf{E}_{\parallel}=\mathrm{Z}_{s}\cdot\left(\hat{\mathbf{n}}\times\mathbf{H}_{\parallel}\right).\$(2)
@@
Given the assumption stated in the introduction, that the metasurface is impenetrable, a positive value for $\langle P_{n}\rangle$ indicates a dissipative response, whereas a negative value indicates amplification, or gain. Given that a lossless metasurface will have $\langle P_{n}\rangle=0$ for all possible incident fields, lossless surfaces have an impedance tensor that is $\rm i$}’ times a Hermitian matrix (a so–called anti–Hermitian operator). In the general case, we can decompose the impedance into Hermitian and anti–Hermitian parts, $Z_{s}=H_{1}+{\rm i}\,H_{2}$, where $H_{1,2}$ are Hermitian matrices. The dissipated power is thus proportional to the Rayleigh quotient of $H_{1}$, implying that passive surfaces (i.e. those without gain) have a positive semi–definite form of $H_{1}$.
@@
III Non–uniqueness of metasurfaces for performing field transformations: imposing reciprocity and passivity
Supposing we want a metasurface to perform a given transformation of the field upon reflection, we can easily find a set of values for the impedance through substituting the desired input and output fields into Eq. (2). To do this, we write our specified complex valued surface fields in terms of their magnitude, e.g. $|E|=(\mathbf{E}_{\parallel}\cdot\mathbf{E}_{\parallel}^{\star})^{1/2}$ and “direction” $\hat{\mathbf{e}}=\mathbf{E}_{\parallel}/|E|$,
$\mathbf{E}_{\parallel}(x,y)$ $=|E(x,y)|\hat{\mathbf{e}}(x,y)$
$\mathbf{H}_{\parallel}(x,y)$ $=|H(x,y)|\hat{\mathbf{h}}(x,y)$ (4)
We note that we can impose any surface fields here, these being the sum of any incident and reflected fields with polarization, phase and amplitude varying arbitrarily across the surface. Substituting this decomposition into the earlier expression for the impedance boundary condition (2) we obtain the requirement
@@
which tells us that at each point on the surface, the impedance tensor must rotate and rescale the complex unit vector $\hat{\mathbf{n}}\times\hat{\mathbf{h}}$ so that it is parallel to the electric field unit vector, $\hat{\mathbf{e}}$. When our desired field transformation has $\hat{\mathbf{e}}$ everywhere parallel to $\hat{\mathbf{n}}\times\hat{\mathbf{h}}_{\parallel}$, this can be achieved with an isotropic metasurface, characterized in terms of a scalar impedance, $\mathrm{Z}_{s}=|E|/|H|$. This is always possible in e.g. a $2D$ geometry, where one field component always lies out of the plane and the other always in the plane.
[…]
$\mathrm{Z}_{s}(x,y)=\frac{|E|}{|H|}\left[\hat{\mathbf{e}}\otimes\left(\hat{\mathbf{n}}\times\hat{\mathbf{h}}\right)^{\star}+\mathbf{v}\otimes\hat{\mathbf{h}}\right$]. (6)
This surface impedance performs the desired arbitrary field transformation (5) for all choices of the complex, position dependent vector, $\mathbf{v}(x,y)$. As just discussed, generally these metasurface parameters are non–reciprocal and non–passive. The remainder of this paper explores the freedom in specifying $\bm{v}$ in Eq. (6), in particular it exploitation to find reciprocal, passive surface parameters.
@@ III.1 Reciprocal metasurfaces:
[…]
$\mathrm{Z}_{s}(x,y)=\frac{|E|}{|H|}\bigg\{e_{1}(x,y)\left[\,\hat{\mathbf{h}}\otimes(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}+\,(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}\otimes\hat{\mathbf{h}}\right]+e_{2}(x,y)\,(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}\otimes(\hat{\mathbf{n}}\times\hat{\mathbf{h}})^{\star}\\ +\alpha(x,y)\,\hat{\mathbf{h}}\otimes\hat{\mathbf{h}}\bigg\}\\$ (7)
@@
From Eq. (LABEL:eq:reciprocal_Z) we can see that, irrespective of the field transformation we seek to enact, through exploiting the non–uniqueness in (6), it is always possible to find a reciprocal metasurface that will perform this transformation. Indeed, we have an infinite family of such surfaces as we are still free to choose the complex valued parameter $\alpha$. Ultimately the physical reason for this design freedom is that at every point we are imposing the response to a given surface field. Of course this leaves the response to the orthogonal surface field unspecified. The somewhat surprising fact is that while the requirement of reciprocity removes some of this redundancy, it is not enough to completely determine the form of the surface impedance. In the next section we will impose the second practical constraint of passivity.
@@ III.2 Passive metasurfaces:
We note that the requirement (9) on the field transformation must be implemented at every point on the surface at the design stage, right from the outset: we cannot hope for a passive surface impedance if our design fields require the metasurface to generate energy! However, even assuming Eq. (9) holds, it may still be that the impedance (LABEL:eq:reciprocal_Z) exhibits gain when probed with other incident fields.
@@
Fig. 2 uses a randomly generated set of surface electric and magnetic fields to show that imposing this constraint indeed results in a passive metasurface. Note that if we impose zero dissipation in the response to the design field, ${\rm Re}[\bar{e}_{2}]=0$, the inequality (12) is generally impossible to fulfill. This shows that reducing the efficiency of the design (i.e. decreasing the reflected power) can help to find a set of metasurface parameters that are, overall passive. We shall return to this point again in the next section.
@@
Figure 2: Randomly generated wave transformations: Using a set of 500500 randomly generated complex unit vectors, $\hat{\mathbf{e}}$ and $\hat{\mathbf{h}}$, representing different choices of wave transformation, we computed the reciprocal surface impedance given by Eq. (LABEL:eq:reciprocal_Z) (with $|E|/|H|=1$, plus the constraint (9)), and the eigenvalues of $(\mathrm{Z}_{s}+\mathrm{Z}_{s}^{\dagger})/2$, which determine the dissipation/gain in the metasurface. (a) $\alpha=0$ for all designs, giving rise to eigenvalues of both signs, generally requiring a non–passive metasurface. (b) $\alpha$ constrained by Eq. (12) (here taking the equality), giving rise to eigenvalues of a positive sign only, showing that all surfaces can be made passive, as desired. (c) Corresponding values of $\alpha$ for the metasurfaces considered in panel (b), showing that in general this parameter must be complex valued.